OpenStudy (anonymous):

I need to write an equation first which i got Y= 1997 -1776. Then i need to solve it thats where i lost i know i need to subtract but i was tring to do the oppistes to both side which makes it adding and then it doesnt work. undo ?? lost and confused The Congress of the United States signed the Declaration of Independence in 1776. Find the difference in years between the year that the Congress signed this document and the year that you were born.

OpenStudy (therealmeeeee):

wrong section please close this and post it in the Math Section

OpenStudy (texaschic101):

The way you wrote the equation, y = 1997 - 1776...is also correct, but in that situation, there is no need to do " the opposite " and add on both sides Now, if you write your equation like this... 1997 - y = 1776 ---- subtract 1997 from both sides 1997 - 1997 -y = 1776 - 1997 -y = - 221-- multiply -1 to both sides to make y positive y = 221 Either equation is correct if you let y be the difference in years.